OpenStudy (anonymous):

A square garden has the dimensions shown: Which of these polynomials could be used to represent the perimeter of the garden? a) 4x^2-3x+9/16 b) 8x+3 c) 16x^2+3x+9/64 d) 4x^2+3x+9/16

OpenStudy (anonymous):



zepdrix (zepdrix):

The perimeter is the distance `around` the garden. So if each side length is \(\large \left(2x+\dfrac{3}{4}\right)\) Then the perimeter will be the sum of all 4 sides, \(\large \left(2x+\dfrac{3}{4}\right)+\left(2x+\dfrac{3}{4}\right)+\left(2x+\dfrac{3}{4}\right)+\left(2x+\dfrac{3}{4}\right)\) Or in other words, \(\large 4\left(2x+\dfrac{3}{4}\right)\)

zepdrix (zepdrix):

Understand how to simplify that?

OpenStudy (anonymous):

prty sure.

OpenStudy (anonymous):

B) 8x+3

zepdrix (zepdrix):

yay, good job \c:/

OpenStudy (anonymous):

tnx